Arundel Company uses aging to estimate uncollectibles. At the end of the fiscal year, December 31, 2018, Accounts Receivable has a balance that consists of:

Dollar Value Age of Account    Estimated Collectible

$215,000 < 30 days old    98.0%

60,000    30 to 60 days old 92.0%

25,000    61 to 120 days old 78.5%

13,000    >120 days old    12.0%

The current unadjusted Allowance for Uncollectible Accounts balance is a debit balance of $2,000 and the Bad Debt Expense accounts has an unadjusted balance of zero. After the adjusting entry is made, what will be the dollar balances in the Allowance for Doubtful Accounts? Round to nearest whole dollar.
Answer #1
Answer a. Date Account Debit Credit
31-Dec Bad debts expense $           27,915
       Allowance for doubtful accounts $       27,915
(Provision made for Bad debts expenses)
Answer b. Balance in allowances for uncollectible $      25,915
Detailed Workings
Calculation of Provision required for Current year
Dollar Value Age of Account Estimated Collectible Estimated Uncollectible % Estimated Uncollectible (Amount)
a b c d = 100-c e = a*d
$215,000 <30 days old 98.00% 2.00% $4,300
60,000 30 to 60 days old 92.00% 8.00% $4,800
25,000 61 to 120 days old 78.50% 21.50% $5,375
13,000 >120 days old 12.00% 88.00% $11,440
Total Estimated Amount of Uncollectibles $25,915
Allowance for Uncollectible Account (Dr) 2000
Bad debt expense for adjusting entry $27,915